OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = [x + (x + \sin ^2 x)^3]^4\]

OpenStudy (lgbasallote):

any hints how to start?

OpenStudy (lgbasallote):

maybe power rule on \((x+\sin^2 x)^3\) first...

OpenStudy (lgbasallote):

\[3(x+\sin^2 x)^2 (1 + 2\sin x \cos x)\]

OpenStudy (zarkon):

multiple chain rules

OpenStudy (lgbasallote):

lol no what i did didnt sound right...i should start from outer

OpenStudy (zarkon):

start with the outside...the power of 4

OpenStudy (lgbasallote):

\[4[x + (x + \sin ^2 x)^3]^3 [1 + 3(x+\sin^2 x)^2 (1 + 2\sin x \cos x)]\]

OpenStudy (lgbasallote):

is that right?

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

ahh! finally